# Optimal Reward Labeling: Bridging Offline Preference and Reward-Based Reinforcement Learning

### v1

@@ 4.2 Theoretical Analysis
Here, we show that many existing PBRL techniques are closely related to some special cases of combining specific offline RL algorithms with our framework under certain conditions.
Offline standard RL algorithms are model-based. First, we consider the case of model-based algorithms. Usually, in the reward-based setting, a model-based offline RL algorithm utilizes the reward signals to learn the reward model of the environment (Yu et al., 2020, 2021). To evaluate the quality of the reward model $\hat{R}$ in predicting the rewards in a dataset $D$, a loss function can be defined as $L=\sum_{(r,s,a)\in D}|\hat{R}(s,a)-r|$. In the case of optimal rewards labels, $r=1$ for the state-actions from the chosen trajectories and $r=-1$ for those from the rejected trajectories. The reward signals influence the update of the reward model through the gradient of the loss. For simplicity, we take one trajectory pair from the dataset as an example. Let $w$ be the training parameters of the reward function (e.g. weights of the neurons in a neural network), the loss and its gradient in the case of optimal reward labeling are given in Eq 1.
$L_{1}$ $=\sum_{(s,a)\in\text{Chosen}}|1-\hat{R}(s,a)|+\sum_{(s,a)\in\text{Rejected}}|\hat{R}(s,a)+1|$ (1)
$=\sum_{(s,a)\in\text{Chosen}}-\hat{R}(s,a)+\sum_{(s,a)\in\text{Rejected}}\hat{R}(s,a)+2\cdot T$
$\frac{\partial L_{1}}{\partial w}$ $=\sum_{(s,a)\in\text{Chosen}}-\frac{\partial\hat{R}(s,a,w)}{\partial w}+\sum_{(s,a)\in\text{Rejected}}\frac{\partial\hat{R}(s,a,w)}{\partial w}$
In recent PBRL works, reward modeling on the preference dataset has also become common. The loss for the reward model in predicting the preference signals and its gradient are given in Eq 2.
$L_{2}=$ $F\Bigl{(}\sum_{(s,a)\in\text{Chosen}}\hat{R}(s,a)-\sum_{(s,a)\in\text{Rejected}}\hat{R}(s,a)\Bigr{)}$ (2)
$\frac{\partial L_{2}}{\partial w}=$ $F^{\prime}\Bigl{(}\sum_{(s,a)\in\text{Chosen}}\hat{R}(s,a)-\sum_{(s,a)\in\text{Rejected}}\hat{R}(s,a)\Bigr{)}$
$\cdot\Bigl{(}\sum_{(s,a)\in\text{Chosen}}\frac{\partial\hat{R}(s,a,w)}{\partial w}-\sum_{(s,a)\in\text{Rejected}}\frac{\partial\hat{R}(s,a,w)}{\partial w}\Bigr{)}$
Note that the loss-link function $F$ is a monotonically decreasing function, so $F^{\prime}(\sum_{(s,a)\in\text{Chosen}}(\hat{R}(s,a))-\sum_{(s,a)\in\text{Rejected}}(\hat{R}(s,a)))$ is always negative.
Comparing the two equations, we find that the gradients for the reward model to update have the same direction in both cases. Even more, if the loss-link function is linear so that $F^{\prime}(\cdot)$ is always a constant, the loss gradients in both cases are equivalent in that they are only different by a scalar multiplier. Therefore, the way how a model-based algorithm utilizes the preference signals under our framework to learn the reward model is the same as that of the current reward-modeling PBRL approaches under the condition that $F$ is a linear function.
@@
In the case where our framework is combined with a reward-based algorithm, we find a similar result. One can adapt a standard model-free offline RL algorithm to the PBRL setting with some intuitive techniques as in Hejna and Sadigh (2024). We show that such methods utilize the preference labels in the same way as our framework when combined with the same model-free algorithm under the condition that the link-loss function $F$ is linear. The detailed analysis is in the Appendix.
@@ 5.1 Experiments Setup
For the baseline methods, to the best of our knowledge, no existing empirical study works in exactly the offline PBRL setting considered in our work. The works that consider our setting are either theoretical studies with no empirical implementation (Zhan et al., 2023; Zhu et al., 2023) or empirical studies focusing on fine-tuning LLM (Ouyang et al., 2022; Rafailov et al., 2024) that cannot be applied to general RL settings. We adopt the IPL algorithm from Hejna and Sadigh (2024) as the existing SOTA method among related works. Different from our method, the IPL algorithm requires access to a preference dataset and a behavior dataset. To ensure that the efficiency of IPL is not underestimated, we allow the algorithm to work with the same preference dataset as our method uses and the whole original D4RL dataset with no reward labels as the behavior dataset, which is strictly more information than our method uses. Li et al. (2024) even point out that one can gain high learning efficiency from access to only the D4RL behavior dataset. In addition, we choose the basic reward modeling method (RM) as a natural baseline. The method first learns a reward model from the preference dataset and then labels the dataset with the reward model. Then, any standard RL algorithms can be applied afterward. This method is similar to the standard pipeline in PBRL Ouyang et al. (2022), where the first step is reward modeling, and the second step is RL.
@@
We choose the popular oracle widely considered in PBRL studies (Hejna and Sadigh, 2024; Kim et al., 2023) where a standard offline RL algorithm is used to train on the dataset with true rewards from the RL environment. The dataset contains the same state actions as the preference-based dataset.
5.2 Learning Efficiency Evaluation
@@
To compare the performance of different learning methods, we show the full training log with different learning methods on different datasets. The training log represents the standard D4RL score of the learned policy by a learning method at each training epoch. The results in Fig 2 show that our method is almost always better than other baseline methods. Note that IPL has additional access to the whole D4RL behavior dataset, which is an advantage compared to our method. In many cases, the learning result of our method can compete with the oracle that trains on the dataset with true reward. We believe our method can sometimes compete with the oracle because the reward dataset already contains more than enough information. As a result, even if the preference dataset contains relatively less information, it is still enough for an efficient learner to find a high-performing policy. Our method can almost always beat the RM baseline because our optimal rewards contain more accurate information than the reward model’s output with respect to the preference signals.
@@ 5.3 Ablation Study
Different size of preference dataset: Here, we examine the efficiency of different methods utilizing the preference signals. For this purpose, we run the algorithms on preference datasets of different sizes. If an algorithm utilizes the preference signals efficiently, then its performance can increase significantly as the number of preference signals increases. We observe in Fig 3 that the learning efficiency of our method increases significantly as the number of preferences increases. In comparison, the learning efficiency of IPL and RM are almost the same for datasets of very different numbers of preference signals. The oracle also has similar performance on the HalfCheetah medium-expert dataset of different sizes. This is likely due to the fact that it always achieves maximal learning efficiency on the dataset. We show the statistical analysis of the ablation studies and additional experiments in the Appendix.
@@
Combing ORL with different standard Offline RL Algorithms. Here, we examine what is the efficiency of our method when combined with different standard offline RL algorithms. We observe in Fig 4 that the learning efficiency is higher if one combines ORL with CQL when learning on the HalfCheetah medium-expert dataset.
5.4 Multiple Preference Labels for Each Pair of Trajectories
Here, we extend our method to the setting where multiple preference labels are given for each pair of trajectories. In this case, the preference labels represent an empirical probability of one trajectory being preferred over the other one. Correspondingly, the loss for predicting the probability given reward labels is given by: $|\bar{p}-f(\sum_{t}r^{i}_{\sigma^{i},t}-\sum_{t}r^{i}_{\bar{\sigma}^{i},t})|$ where $\bar{p}$ is the empirical probability. We add a regularization term to the loss to ensure the uniqueness of the optimal reward labels. In this case, finding the optimal rewards requires the knowledge of the link function from rewards to preferences. In experiments, we construct the dataset using the same process as before, except that we generate $10$ preference labels for each pair of trajectories. We also compare the difference in learning efficiency between this setting and the previous single preference setting. The results in Fig 5 show that our algorithm is equally efficient in both settings.
Figure 4: Comparison between the learning efficiency of ORL combined with different standard offline RL algorithms.
Figure 5: Comparison between the cases where single or multiple preference labels are given to each pair of trajectories.
@@ 6 Conclusion and Limitation
In this work, we propose a framework ORL that bridges the gap between offline PBRL and standard RL. We show that one can easily achieve high learning efficiency on PBRL problems by combining ORL with any efficient standard offline RL algorithm. Our framework is limited to the typical offline learning setting and does not answer the question of which trajectories are more worthy of receiving preference labels. Our experimental evaluation is limited to continuous control problems, the standard benchmark in RL studies. The next step can be extending our work to other PBRL settings and other tasks such as language model finetuning.
Appendix A Proof for Lemma 4.1

### v3

@@ 4.2 Theoretical Analysis
In this section, we show that existing PBRL techniques are closely related to some special cases of combining specific offline RL algorithms with our framework. Under certain conditions, they can even be equivalent.
Offline standard RL algorithms are model-based. First, we consider the case of model-based algorithms. Usually, a model-based offline RL algorithm utilizes the reward signals to learn the reward model of the environment (Yu et al., 2020; 2021). Formally, we characterize the reward modeling process in a general model-based offline RL algorithm in Definition 4.3.
Definition 4.3.
(reward modeling in model-based approaches) Given a reward-based dataset $\mathcal{D}_{\mathcal{R}}=\{(s_{i},a_{i},r_{i}),i\in[N]\}$, the reward modeling process is solving the optimization problem
$\min_{\widehat{\mathcal{R}}}\sum_{(r,s,a)\in D}|\widehat{\mathcal{R}}(s,a)-r|$
, where ^$\widehat{\mathcal{R}}:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]$ is a reward model.
Here, we extend Definition 4.3 to the case of the preference dataset with binary reward labeling. Given a preference dataset $\mathcal{D}=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$, after applying the binary reward labeling through Alg 1, the reward-based dataset $\mathcal{D}_{\mathcal{R}}$ consists of all state-actions from $\mathcal{D}$. For each state-action $(s,a)\in\mathcal{D}_{\mathcal{R}}$, the reward label is $+1$ if the state-action comes from a chosen trajectory and $-1$ otherwise. In this case, the optimization problem becomes
$\min_{\widehat{\mathcal{R}}}\sum_{i\in[N]}\sum_{(s,a)\in\tau^{i}_{1}}|1-\widehat{\mathcal{R}}(s,a)|+\sum_{(s,a)\in\tau^{i}_{2}}|\widehat{\mathcal{R}}(s,a)+1|:=\sum_{i\in[N]}\min_{\widehat{\mathcal{R}}}\mathcal{L}_{1}(\tau^{i}_{1},\tau^{i}_{2},\mathcal{R}).$
Next, we formally introduce the reward modeling process based on the preference dataset directly in Definition 4.4. As we explained earlier, this process is similar to solving the optimal reward label problem in Definition 4.1 and standard in current PBRL studies.
Definition 4.4.
(reward modeling on preference signals) Given a preference-based dataset $\mathcal{D}=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$, the reward modeling process is solving the optimization problem below:
$\min_{\widehat{\mathcal{R}}}\sum_{i\in[N]}F\Bigl{(}\sum_{(s,a)\in\tau^{i}_{1}}\hat{R}(s,a)-\sum_{(s,a)\in\tau^{i}_{2}}\hat{R}(s,a)\Bigr{)}:=\sum_{i\in[N]}\min_{\widehat{\mathcal{R}}}\mathcal{L}_{2}(\tau^{i}_{1},\tau^{i}_{2},\mathcal{R}).$
Finally, in Theorem 4.5, we formally show the connection between the two methods.
Theorem 4.5.
Given a preference dataset $\mathcal{D}=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$, reward modeling can be performed either on the preference dataset directly as Definition 4.4 or on the reward dataset as Definition 4.3 with the dataset generated binary reward labeling through Alg 1. The two methods are connected in the following three cases:
  1. 1.
    When there is no overlap between the trajectories in the dataset, the optimal solutions in both methods are the same: $\arg\min_{\widehat{\mathcal{R}}}\sum_{i\in[N]}\mathcal{L}_{1}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})=\arg\min_{\widehat{\mathcal{R}}}\sum_{i\in[N]}\mathcal{L}_{2}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})$
  2. 2.
    If the link-loss function $\mathcal{F}$ is linear, then the optimization problems in both methods are equivalent: $\sum_{i\in[N]}\mathcal{L}_{1}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})=C_{1}\cdot\sum_{i\in[N]}\mathcal{L}_{2}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})+C_{2},$ where $C_{1},C_{2}$ are constant scalars.
  3. 3.
    Let $w$ be the parameter of the reward function $\mathcal{R}$. For each trajectory pair, the gradients of its contribution to the optimization goal on the reward function parameter have the same direction in the two methods: $\frac{\partial\mathcal{L}_{1}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})}{\partial w}/\|\frac{\partial\mathcal{L}_{1}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})}{\partial w}\|=\frac{\partial\mathcal{L}_{2}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})}{\partial w}/\|\frac{\partial\mathcal{L}_{2}(\tau^{i}_{1},\tau^{i}_{2},\widehat{\mathcal{R}})}{\partial w}\|$.
The proof for Theorem 4.5 can be found in the appendix. The first case shows that in the most practical scenario where there is no overlap between trajectories, the reward modeling in both methods approximates the same optimal reward function. The second case shows that the reward modeling in both methods is equivalent if the link-loss function is linear. The third case shows that the reward model update based on each trajectory pair in both methods is the same. In conclusion, the way how a model-based algorithm utilizes the preference signals for reward modeling under our framework is closely related to that of the current reward-modeling PBRL approaches.
@@
In the case where our framework is combined with a reward-based algorithm, we find a similar result. One can adapt a standard model-free offline RL algorithm to the PBRL setting with some intuitive techniques as in Hejna & Sadigh (2024). We show that such methods utilize the preference labels in the same way as our framework when combined with the same model-free algorithm under the condition that the link-loss function $F$ is linear. The detailed analysis is in the Appendix.
@@ 5.1 Experiments Setup
For the baseline methods, to the best of our knowledge, no existing empirical study works in exactly the standard offline PBRL setting considered in our work. The works that consider our setting are either theoretical studies with no empirical implementation (Zhan et al., 2023; Zhu et al., 2023) or empirical studies focusing on fine-tuning LLM (Ouyang et al., 2022; Rafailov et al., 2024) that cannot be applied to general RL settings. We adopt the IPL algorithm from Hejna & Sadigh (2024) as the existing SOTA method among related works. Different from our method, the IPL algorithm requires access to a preference dataset and a behavior dataset. To ensure that the efficiency of IPL is not underestimated, we allow the algorithm to work with the same preference dataset as our method uses and the whole original D4RL dataset with no reward labels as the behavior dataset, which is strictly more information than our method uses. In addition, we choose the basic reward modeling method (RM) as a natural baseline. The method first learns a reward model from the preference dataset and then labels the dataset with the reward model. Next, any standard RL algorithms can be applied afterward. This method is similar to the standard pipeline in PBRL Ouyang et al. (2022), where the first step is reward modeling, and the second step is RL.
@@
We choose the popular oracle widely considered in PBRL studies (Hejna & Sadigh, 2024; Kim et al., 2023) where a standard offline RL algorithm is used to train on the dataset with true rewards from the RL environment. The dataset contains the same state actions as the preference-based dataset.
5.2 Learning Efficiency Evaluation without Trajectory Overlap
@@
Here, we study the most practical case where there is no overlap between trajectories, and each state-action is unique. To straightforwardly compare the performance of different learning methods, we represent the learning efficiency of an algorithm by the performance of the policy learned at the last epoch. In the Appendix, we show the full training log with different learning methods on different datasets. The scores in Table 1 is the standard D4RL score of the learned policy. We observe that our method is better than other baseline methods in general. Note that IPL has additional access to the whole D4RL behavior dataset, which is an advantage compared to our method. In many cases, the learning result of our method can compete with the oracle that trains on the dataset with true reward. We believe our method can sometimes compete with the oracle because the reward dataset already contains more than enough information. As a result, even if the preference dataset contains relatively less information, it is enough for an efficient learner to find a high-performing policy. Our method is, in general, more efficient than the RM baseline. We believe the reason is that our optimal rewards contain more accurate information than the reward model’s output with respect to the preference signals.
Oracle BRL RM IPL
HalfCheetah M 98.05 $\pm$ 0.24 75.93 $\pm$ 3.64 56.52 $\pm$ 0.74 39.71 $\pm$ 2.17
HalfCheetah MR 62.61 $\pm$ 8.70 61.72 $\pm$ 0.60 57.36 $\pm$ 1.60 21.14 $\pm$ 1.14
HalfCheetah ME 87.12 $\pm$ 11.39 92.17 $\pm$ 1.67 87.52 $\pm$ 3.98 34.91 $\pm$ 0.38
Hopper M 57.16 $\pm$ 1.25 36.39 $\pm$ 3.90 43.38 $\pm$ 1.00 51.15 $\pm$ 11.89
Hopper MR 94.07 $\pm$ 6.17 28.5 $\pm$ 10.29 24.44 $\pm$ 3.80 7.48 $\pm$ 0.42
Hopper ME 97.71 $\pm$ 16.99 97.57 $\pm$ 15.08 74.91 $\pm$ 13.68 28.44 $\pm$ 15.56
Walker2d M 80.77 $\pm$ 4.72 49.7 $\pm$ 11.06 58.55 $\pm$ 4.46 67.33 $\pm$ 7.66
Walker2d MR 75.44 $\pm$ 3.84 34.94 $\pm$ 12.21 13.26 $\pm$ 9.11 14.87 $\pm$ 3.97
Walker2d ME 111.74 $\pm$ 0.13 109.94 $\pm$ 1.76 90.1 $\pm$ 35.01 85.76 $\pm$ 10.43
Sum Totals: 764.67 586.86 506.04 350.79
Table 1: Performance of different learning algorithms on the dataset without overlapped trajectories. ‘M’ represents the medium dataset, ‘MR’ represents the medium replay dataset, and ‘ME’ represents the medium expert dataset.
5.3 Overlapped Trajectories
Here, we empirically investigate the case where the trajectory clips in the dataset share the same state-actions. Recall that when there is overlap between trajectories (state-actions in the dataset are not unique), the binary labeling strategy is no longer optimal. The reward modeling is approximating the optimal labels, which can be a stronger baseline in this case. Therefore, in this subsection, we focus on comparing which reward labeling method can give more informative rewards leading to higher learning efficiency. Specifically, we consider two typical structures of overlap between trajectories.
I: Same trajectory compared to multiple different trajectories. In this case, we study a more practical scenario of overlapped trajectories: the same trajectory clip is compared with multiple trajectory clips. More specifically, given a pool of trajectory clips, we sample a portion of them and compare it with multiple different clips from the pool. We set the portion of number of comparisons from low to high to cover a wide range of the degree of trajectory overlap.
Figure 2: Training log of learning with a method on datasets with overlapped trajectories. The percentage is the portion of trajectories clips that are compared for multiple times compared to all clips. The multiplier is the number of times of multiple comparison. To understand the degree of overlap, in the case of $20\%\times 4$, $80\%$ of the trajectories pairs have a trajectory clip that is compared for multiple times.
The results in Figure 2 show that the BRL method is more efficient than the reward modeling. To understand the reason behind, we check the gap between the reward labels given by BRL and reward modeling. We find that the gap is much larger in the BRL method compared to the reward modeling method, indicating that the binary reward labels keep more information during the feedback signal transition from preference to scalar rewards. The exact comparison results can be found in the Appendix. This could be the reason why BRL method is more efficient.
II: The same trajectory pair compared multiple times. Here, we investigate the scenario where multiple preference signals are given to the same pair of trajectory clips. This could happen if multiple users are asked to provide preferences on the same trajectory pairs, and different users may have different judgments. In this case, the preference labels represent an empirical probability of one trajectory being preferred over the other one. Correspondingly, the loss for predicting the probability given reward labels is given by: $|\bar{p}-f(\sum_{t}r^{i}_{\sigma^{i},t}-\sum_{t}r^{i}_{\bar{\sigma}^{i},t})|$ where $\bar{p}$ is the empirical probability. We add a regularization term to the loss to ensure the uniqueness of the optimal reward labels. In this case, finding the optimal rewards requires the knowledge of the link function from rewards to preferences. In experiments, we construct the dataset using the same process as before, except that we generate $10$ preference labels for each pair of trajectories.
Figure 3: Training log of learning with a method on datasets where $10$ preference labels are given to each trajectory pair.
The results in Figure 3 show that in the two datasets we test with, the BRL method is more efficient or as efficient as the reward modeling method. In conclusion, compared to the standard reward modeling method, with or without overlaps between trajectories, the BRL method generally performs better than the reward modeling method. In addition, the BRL method requires no training for function approximation, which is more computationally efficient.
@@ 5.4 Ablation Study
Different size of preference dataset: Here, we examine the efficiency of different methods utilizing the preference signals. For this purpose, we run the algorithms on preference datasets of different sizes. If an algorithm utilizes the preference signals efficiently, then its performance can increase significantly as the number of preference signals increases. We observe in Fig 4 that except for IPL, the learning efficiency of other methods increases significantly as the number of preferences increases. The performance of our BRL method degrades slower than that of the RM method as the dataset’s size decreases, suggesting that it utilizes the data more efficiently. When the dataset is large enough, the BRL method can perform as well as the oracle. The reason could be that the information in the dataset is abundant enough that even if a lot of information is lost from true reward to preference feedback, an agent is still able to recover the optimal policy from the preference labels.
@@
In the appendix, we show the learning efficiency of combining BRL with different standard Offline RL Algorithms.
@@ 6 Conclusion and Limitation
In this work, we propose a framework BRL that bridges the gap between offline PBRL and standard RL. We show that one can easily achieve high learning efficiency on PBRL problems by combining BRL with any efficient standard offline RL algorithm. Our framework is limited to the typical offline learning setting and does not answer the question of which trajectories are more worthy of receiving preference labels. Our experimental evaluation is limited to continuous control problems, the standard benchmark in RL studies.
7 Reproducibility
In the main paper, we explain the setting of the problem we study. The proofs for all theorems and lemmas can be found in the appendix. The codes we use for the experiments can be found in the supplementary materials.
Appendix A Combining BRL with Model-Free Standard Offline RL Algortihms
It is typical for a model-free RL algorithm to learn the Q function of the environment (Fujimoto & Gu, 2021; Cheng et al., 2022). The Q function represents the long-term rewards for choosing an action at a state. In the reward-based setting, the Q function is learned with ‘Bellman Loss’ based on the reward signals. Bellman loss acts as a criterion for the quality of the $Q$ function. Given a tuple $(s,a,r,s^{\prime})$, the Bellman loss on the data tuple is defined as $|Q(s,a)-(r+\gamma\cdot\max_{a^{\prime}}Q(s^{\prime},a^{\prime})|$. Formally, in Definition A.1, we characterize the Q-learning process in a typical model-free method.
Definition A.1.
(Q-learning on reward signals in a general model-free RL method with binary reward labels) Given a preference dataset $\mathcal{D}=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$. Let $\mathcal{D}_{\mathcal{R}}$ be the corresponding reward dataset with binary reward labels given by Alg 1. The Q-learning process for a general model-free RL method is solving the optimization problem below:
$\min_{\hat{Q}}\sum_{i\in[N]}\Bigl{(}\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}|\hat{Q}(s,a)-(1+\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))|+\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}|\hat{Q}(s,a)-(-1+\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a^{\prime}))|\Bigr{)}.$
For simplicity, we denote the optimization goal as $\min_{\hat{Q}}\sum_{i\in[N]}\mathcal{L}_{3}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q}).$
If one wants to modify a model-free reward-based algorithm to work with preference signals directly, a straightforward way is to replace the Bellman error with the prediction error of the $Q$ function on the preference signals. Based on this method, Hejna & Sadigh (2024) developed the IPL algorithm for PBRL by modifying the IQL algorithm (Kostrikov et al., 2021), a famous method for the standard RL problem. Specifically, the underlying reward function for a $Q$ function is given by $\hat{r}(s,a)=\hat{Q}(s,a)-\gamma\cdot\max\hat{R}(s^{\prime},a^{\prime})$. This underlying reward function can be used to predict the probability that the chosen trajectory is preferred to represent the prediction of the $Q$ function. To reflect that the reward function is bounded, we require $Q(s,a)-\gamma\cdot Q(s^{\prime},a)\in[-1,1]$ for all $(s,a,s^{\prime})\sim\mathcal{D}$ so that the corresponding reward function of $Q$ is bounded. Formally, in Definition A.2 we show how a method to perform Q-learning directly on preference signals.
Definition A.2.
(Q-learning on preference signals through a modified Bellman-Loss) Given a preference dataset $\mathcal{D}=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$, the Q-learning on preference signals through a modified Bellman-Loss is solving the optimization problem below:
$\min_{\hat{Q}}\sum_{i\in[N]}F\Bigl{(}\sum_{(s,a,s^{\prime})\in\tau^{i}_{1}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a)))-\sum_{(s,a,s^{\prime})\in\tau^{i}_{2}}(\hat{Q}(s,a)-\gamma\cdot\max_{a^{\prime}}\hat{Q}(s^{\prime},a))\Bigr{)}.$
For simplicity, we denote the optimization goal as $\min_{\hat{Q}}\sum_{i\in[N]}\mathcal{L}_{4}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q}).$
The two methods are closely connected in the optimization problem they solve, and we formally show their connection in three cases in Theorem A.3.
Theorem A.3.
Given a preference dataset $\mathcal{D}=\{(\tau_{1}^{i}\succ\tau_{2}^{i})\},i\in[N]$, q-learning can be performed either on the preference dataset directly as Definition A.2 or on the reward dataset as Definition A.1 with the dataset generated binary reward labeling through Alg 1. The two methods are connected in the following three cases:
  1. 1.
    When there is no overlap between the trajectories in the dataset, the optimal solutions in both methods are the same: $\arg\min_{\hat{Q}}\sum_{i\in[N]}\mathcal{L}_{3}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})=\arg\min_{\hat{Q}}\sum_{i\in[N]}\mathcal{L}_{4}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})$
  2. 2.
    If the link-loss function $\mathcal{F}$ is linear, then the optimization problems in both methods are equivalent: $\sum_{i\in[N]}\mathcal{L}_{3}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})=C_{1}\cdot\sum_{i\in[N]}\mathcal{L}_{4}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})+C_{2},$ where $C_{1},C_{2}$ are constant scalars.
  3. 3.
    Let $w$ be the parameter of the reward function $\mathcal{R}$. For each trajectory pair, the gradients of its contribution to the optimization goal on the reward function parameter have the same direction in the two methods: $\frac{\partial\mathcal{L}_{3}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})}{\partial w}/\|\frac{\partial\mathcal{L}_{3}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})}{\partial w}\|=\frac{\partial\mathcal{L}_{4}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})}{\partial w}/\|\frac{\partial\mathcal{L}_{4}(\tau^{i}_{1},\tau^{i}_{2},\hat{Q})}{\partial w}\|$.
Appendix B Proof
B.1 Proof for Lemma 4.2